The demand curve facing a monopoly is P= 100 +5Q. The firm cost is c(Q)=10+5Q. What is profit maximising output?



1
Expert's answer
2021-07-07T09:01:34-0400

Given;

"P= 100 +5Q."

"c(Q)=10+5Q."


"TR=P\times Q\\=(100+5Q)Q\\=100Q+5Q^2\\MR=100+10Q"


"c(Q)=10+5Q\\MC=5"


The profit is maximized by producing at a point where MC=MR


"5=100+10Q\\-95=10Q\\Q=-9.5"